OpenStudy (anonymous):

Describe the possible solution sets for a system of two linear equations, and explain why these are the only possibilities.

OpenStudy (anonymous):

is there an example or no?

OpenStudy (anonymous):

like do we have the 2 linear equations?

OpenStudy (anonymous):

or do we come up with the linear equations?

OpenStudy (anonymous):

no ithis is th who question

OpenStudy (anonymous):

you could do x+y=10 2x+2y=20 and say it is consistent

OpenStudy (anonymous):

because when you divide the second equation by 2 you get x+y=10 which doesn't contradict the first statement